Calculate the resting potential of a huge axon of a squid if concentration of sodium ions outside is 480 mMol and inside is 42 mMol (t=30C).
1
Expert's answer
2020-05-29T10:03:20-0400

The resting potential can be found by the Nernst equation:


"\phi=\frac{RT}{zF}\text{ ln}\frac{[\text{Na}^+_\text{outside}]}{[\text{Na}^+_\text{inside}]}=\\
\space\\
=\frac{8.314\cdot(273+30)}{1\cdot 9.65\cdot10^4}\text{ ln}\frac{480}{42}=0.0636\text{ V},"

or 63.6 mV.